OpenStudy (anonymous):

13/4*4/52 mulitiply in simpliest form

OpenStudy (anonymous):

13/4 + 4/3 = (39+16)/12 = 55/12 =4 7/12 in mixed frac i thought it ws + nywaz ....

OpenStudy (anonymous):

it's 1/4

Parth (parthkohli):

\[\LARGE {13 \over 4} * {4 \over 52} = {1 \over 1} * {1 \over 4} = {1 \over 4} \]

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

i beat you to it

OpenStudy (anonymous):

:P

Parth (parthkohli):

Tai, you didn't explain

OpenStudy (anonymous):

calculator ftw - medal please :]

Parth (parthkohli):

@Rohangrr It's a multiply sign lol

OpenStudy (anonymous):

oops!

OpenStudy (anonymous):

dizzy hint... \[\LARGE {13\over 4}\cdot {4\over 52}\] \[\LARGE {13\over \cancel{4}}\cdot {\cancel{4}\over 52}\] ... @Tai Openstudy is about teaching, not just giving blindly answers. Your answer is worthless if you don't explain it! ;) lol

Parth (parthkohli):

@Kreshnik how did you put the cut thing?

OpenStudy (anonymous):

\cancel { 4 }

Parth (parthkohli):

Cool bro.. LATEX expert :P

Parth (parthkohli):

\[\LARGE {\cancel {4} \over 3}\] Yeah

OpenStudy (anonymous):

show off...

OpenStudy (anonymous):

:p